You are offered ​$80,000 today or ​$340,000 in 15 years. Assuming that you can earn 11 percent on your​ money, which should you​ choose?
Answer #1

Present value of money received in 15 years = future value ÷( 1+r)^15

Present value of option 1 : $340000÷{(1+0.11)^15}

Present value of option 1 = $71,061.48

Present value of option 2 : $80000

We should choose option 2 of receive 80000 today.
Answer #2

SOLUTION :


It requires to compare the present values of options to select the better one. Higher the PV, better is the option. 


Accordingly, we determine PVs of the two options .


Option 1 :


$80000 today .

Its PV = $80000 .


Option 2 :


Receiving $340000 in 15 years from now. 

Required rate, r = 11% = 0.11

=> 1 + r = 1.11


PV of Option 2 

= FV/(1 + r)^n 

= 340000 / 1.11^15 

= 71061.48 ($) 


PV of Option 2 is = $71061.48 (ANSWER)


So, PV of Option 1 is higher than that of Option 2. 

Hence, Option 1 should be accepted. (ANSWER).



answered by: Tulsiram Garg